OpenStudy (anonymous):

Use l’Hopital’s rule to show that f is continuous at x=-1.

OpenStudy (anonymous):

I worked out that the limit at x=-1 is \[\frac{ \ln(b) }{ \ln(a) }\]. but this is only near x=-1 not at x=-1

OpenStudy (primeralph):

differentiate the first piece top and bottom, plug in -1.

OpenStudy (anonymous):

ye, i did that. my question is its the limit as x approaches -1. l’Hopital’s will give the derivative as x approches -1. However, this says nothing about the point AT x = -1. if might not be defined, thats the question